A football player pushes a football sled with an applied a horizontal force of 103 N for a distance of 15 m. If this exercise took 25 s to complete, calculate the power produced by the football player. 


1
Expert's answer
2021-01-25T08:28:50-0500

Answer

the power produced by the football player is given by

"P=\frac{F\times d}{t}\\
 =\frac{103\times 15}{25}=61.8Watts"